OpenStudy (anonymous):

The diameter of a cylindrical propane gas tank is 4 feet. The total volume of the tank is 527.8 cubic feet. Find the length of the tank. (Round your answer to the nearest foot.)

OpenStudy (anonymous):

the equation for a cylinder is v=h(pi)r^2 527.8=h(pi)(4)^2 solve for h